The base of a prism is a regular pentagon with one side measuring 15 cm. Its

Altitude is 32 cm. Find the following;

  1. Lateral surface Area in cm2.
  2. The total surface area in cm2.
  3. The volume of the prism in cm3.
1
Expert's answer
2022-03-29T12:18:06-0400

Let "AB=BC=CD=DE=AE" then "A"1"B"1"=B"1"C"1"=C"1"D"1"=D"1"E"1 "=E"1"A"1

Then the area of the upper and lower bases of the prism: "S\scriptsize{base} \normalsize =\frac{AB^2\sqrt{25+10\sqrt{5}}}4=\frac{225\sqrt{25+10\sqrt{5}}}4",

both bases: "S^*=2S=\frac{225\sqrt{25+10\sqrt{5}}}2"

1.

A side face has the area "S=AB\cdot AA"1 , all five side faces have the area "S^{**}=5S=5\cdot15\cdot32=2400" cm2

2.

The total surface area: "S=S^*+S^{**}=\frac{225\sqrt{25+10\sqrt{5}}}2+2400\thickapprox3175,2148" cm2

3.

The volume of the prism: "V=S\scriptsize{base}\normalsize \cdot AA"1 "=\frac{225\sqrt{25+10\sqrt{5}}}4\cdot32=1800\sqrt{25+10\sqrt{5}}" "\thickapprox12387,4373" cm3